OpenStudy (anonymous):

Complete the square to solve the equation: x^2+6x=-7

OpenStudy (anonymous):

what is half of \(6\) ?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

ok good then write \[(x+3)^2=-7+?\] now what is \(3^2\) ?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

right so it is \[(x+3)^2=-7+9\] or \[(x+3)^2=2\]

OpenStudy (anonymous):

take the square root of both sides, get \[x+3=\pm\sqrt2\] then subtract \(3\)to finish with \[x=-3\pm\sqrt2\]

OpenStudy (anonymous):

Thank you